The end of a diving board undergoes simple harmonic motion with the velocity shown below: vy(t)=vmax*sin((2*pi*t)/T) where vmax = 4.64 m/s and T = 0.553 seconds. (calculator must be in radians)

What is the angular frequency of the board's velocity?

What is the amplitude of the board's position?
Answer #1

Velocity is given by:

Vy(t) = Vmax*sin (2*pi*t/T)

Now given that Vmax = 4.64 m/sec

T = time period = 0.553 sec

Part A.

Angular frequency is given by:

w = 2*pi/T

w = 2*pi/0.553

w = 11.36 rad/sec

Part B.

Relation between amplitude and max velocity is given by:

Vmax = A*w

A = Vmax/w

A = 4.64/11.36

A = 0.41 m = Amplitude

Please Upvote.